# Resolving Gauge Ambiguities of the Berry Connection in Non-Hermitian Systems

### v1

@@ Abstract
Non-Hermitian systems display spectral and topological phenomena absent in Hermitian physics; yet, their geometric characterization can be hindered by an intrinsic ambiguity rooted in the eigenspace of non-Hermitian Hamiltonians, which becomes especially pronounced in the pure quantum regime. Because left and right eigenvectors are not related by conjugation, their norms are not fixed, giving rise to a biorthogonal ${\rm GL}(N,{\mathbb{C}})$ gauge freedom. Consequently, the standard Berry connection admits four inequivalent definitions depending on how left and right eigenvectors are paired, giving rise to distinct Berry phases and generally complex-valued holonomies. Here we show that these ambiguities and the emergence of complex phases are fully resolved by introducing a covariant-derivative formalism built from the metric tensor of the Hilbert space of the underlying non-Hermitian Hamiltonian. The resulting unique Berry connection remains real-valued under an arbitrary ${\rm GL}(N,{\mathbb{C}})$ frame change, and transforms as an affine gauge potential, while reducing to the conventional Berry (or Wilczek–Zee) connection in the Hermitian limit. This establishes an unambiguous and gauge-consistent geometric framework for Berry phases, non-Abelian holonomies, and topological invariants in quantum systems described by non-Hermitian Hamiltonians.
@@ I Introduction
Non-Hermitian Hamiltonians naturally emerge in a wide variety of experimental platforms, including photonic systems with engineered gain and loss [29, 11, 28], mechanical and electrical metamaterials with active or nonreciprocal elements [13, 32], acoustic lattices [43, 18], cold-atom systems with controlled dissipation [39, 4], circuit-QED and microwave platforms [30], as well as synthetic quantum matter  [28, 6] to name a few. In these systems, the nontrivial phenomena such as non-Hermitian skin effects [42, 22, 24] and non-Bloch band topology [42, 27] demonstrate that their geometric properties encode unique features not encountered in Hermitian counterparts.
A defining feature of the non-Hermitian Hamiltonians is the biorthogonal structure of their eigenvectors: right and left eigenstates are independent, and need not be related by Hermitian conjugation [25]. As a consequence, each eigenmode retains an intrinsic ‘normalization’ gauge freedom, namely,
@@
whose transformations under $c_{n}(\boldsymbol{\lambda})$ differ because left and right vectors transform oppositely [12, 34, 38]. As a result, these four connections are generically not gauge equivalent. Moreover, for any parameter-dependent scaling with $|c_{n}|\neq 1$, all ${\cal A}^{\alpha\beta}$ acquire different gauge-dependent imaginary components, rendering the associated Berry phases complex and curvatures non-unique [34]. Specifically, under such rescaling in Eq. (1) with $c_{n}=r_{n}e^{i\phi_{n}}$, the left-right Berry connection for a single-band transforms as
@@
Clearly, for arbitrary real-valued scaling factor $r_{n}\neq 1$, the Berry connection becomes complex-valued even if the original connection ${\cal A}^{LR}$ is purely real. This implies that a state undergoing adiabatic evolution may, in general, exhibit geometric exponential damping or amplification along open-path trajectories, irrespective of the specific form of the Berry connection ${\cal A}^{LR}$. On the other hand, as Eq. (3) suggests, for closed loops the geometric phase is insensitive to such rescaling-induced terms $r_{n}$ and depends solely on the intrinsic properties of ${\cal A}^{LR}$, provided that the $\phi_{n}$ is single-valued. Furthermore, when the Berry connection in Eq. (2) is complex, the accumulated Berry phase along a closed loop is generally complex as well.
While such complex geometric phases pose no conceptual difficulty in classical and semiclassical dissipative systems, where the evolving state is a physical field that may naturally attenuate or amplify along a given trajectory in parameter space [35], and where such geometric effects can even be observed and exploited, as recently demonstrated in Refs. [38, 23], the situation is fundamentally different in the quantum domain. When the wave function encodes probability amplitudes of the system’s possible configurations, rather than a physical field, its norm represents the total probability and must therefore remain fixed under any admissible evolution, whether generated by a Hermitian or a non-Hermitian operator [10]. Although it is common to enforce norm preservation through ad hoc renormalization of the quantum state, this procedure raises concerns about the consistency of such an approach in genuine non-Hermitian quantum settings [36].
Consequently, this invokes a natural question: Can one define a unique and real-valued Berry connection on the eigenspace of non-Hermitian Hamiltonians that would also remain real under arbitrary ${\rm GL}(N,\mathbb{C})$ frame transformations, while still retaining the essential properties of a gauge potential, possibly in some modified form? Such a connection, which ensures the geometric norm preservation of the adiabatically evolving states, would lead to a uniquely defined real-valued phases, and a consistent geometric and topological structure on the eigenvector bundle of the quantum non-Hermitian evolution operators. In this work we construct such a Berry connection and show that it naturally emerges from the metric associated with the ‘deformed’ non-Hermitian Hamiltonian Hilbert space.
@@ II Metric tensor and covariant derivative in the Hilbert space of non-Hermitian Hamiltonians
corresponding to some in general time-dependent Hermitian Hamiltonian $H^{H}=SHS^{-1}+i(\partial_{t}S)S^{-1}$ [20]. The latter expression coincides with the gauge transformation law for the Hamiltonian viewed as a gauge potential, with the transformation rendering it Hermitian. Indeed, combining together Eqs. (5)–(7) one immediately obtains $\langle\phi^{H}|\phi^{H}\rangle=1$, which is a standard definition of the inner product on the Hilbert space of Hermitian Hamiltonians (see also Fig. 1111Note that the Hermitizing map $S$ possesses a unitary gauge freedom and thus is not unique [20]. Indeed, by setting $S\to US$, with some unitary $U$, the metric $\eta$ in Eq. (6) remains invariant. We implicitly assume that such $U$ is always continuous across the system parameter space under consideration, ensuring that the associated covariant derivative is well defined, as discussed later in the text..
@@
where the $S^{-1}\partial_{\lambda}S=\Gamma^{\lambda}$ term serves as the connection component on the $\eta$-deformed Hilbert space and obeys the metric-compatible equation for the connection: $\partial_{\lambda}\eta=\eta\Gamma^{\lambda}+\Gamma^{\lambda\dagger}\eta$  [^2] More specifically, $\partial_{\theta}\langle\psi|\psi\rangle_{\eta}=\langle D_{\theta}\psi|\psi\rangle_{\eta}+\langle\psi|D_{\theta}\psi\rangle_{\eta}$=0, implying $\langle D_{\theta}\psi|\psi\rangle_{\eta}=-\langle\psi|D_{\theta}\psi\rangle_{\eta}$. This condition does not, in general, enforce parallel transport of the state, determined by the expression $D_{\lambda}|\psi\rangle=0$, but merely guarantees norm preservation of a state when it is transported in $\boldsymbol{\lambda}$-space [19]. This situation is fully analogous to the Hermitian case: although the norm of an eigenstate $|\phi^{H}\rangle$ is fixed in the parameter space, $\partial_{\lambda}\langle\phi^{H}|\phi^{H}\rangle=0$, this does not imply that the state itself is constant, i.e., $\partial_{\lambda}|\phi^{H}\rangle=0$..
Furthermore, this connection is curvature-free, as it is generated by a pure-gauge structure (i.e., the Maurer-Cartan form) $S^{-1}dS$, implying $D_{\mu},D_{\nu}]=0 [$26, 19]. Meaning that the Hilbert space, endowed with the metric $\eta$, is intrinsically flat despite the seeming deformation (see Fig. 1).
@@ III Covariant Berry connection
In the Hermitian limit, for a fixed eigenframe, the CBC reduces to the standard Berry connection, since $\eta\to I$ and $D_{\lambda}\to\partial_{\lambda}$. Likewise, when the metric constant, that is $\Gamma^{\lambda}=0$, the CBC is simplified to the usual left-right non-Hermitian Berry connection given in Eq. (2). In this sense, the standard non-Hermitian left–right connection ${\cal A}^{LR}$ is appropriate in the genuine quantum regime only when the Hilbert-space metric is fixed (i.e., $\eta=\rm const$). Moreover, via the similarity transformation in Eq. (7), the CBC can also be written as ${\mathbb{A}}^{\lambda}_{mn}=i\langle\phi^{H}_{m}|\partial_{\lambda}|\phi^{H}_{n}\rangle$, which is the conventional Berry connection for a Hermitian system. Thus the CBC in Eq. (9) is manifestly real-valued.
@@
Here, the distortion tensor $\Xi$ is determined by the difference of the metric connections $\Gamma^{\prime\lambda}-\Gamma^{\lambda}$, originating from the non-unitary (positive-definite) part $P$ in the polar decomposition of $T=UP$. In other words, the distortion compensates the local rescaling of eigenvector norms induced by $P$. In particular, for a purely unitary transformation $T=U$, the metric remains invariant $\eta^{\prime}=\eta$, which makes $(\Gamma^{\prime}-\Gamma)$ a pure gauge  [^3] Indeed, as noted earlier in the text, for the fixed metric $\eta$, the two connections $\Gamma$ and $\Gamma^{\prime}$ may arise from the two Hermitized maps $S$ and $S^{\prime}=US$, respectively, which differ by a smooth unitary $U$. According to the definition of the connection in Eq. (8), these two connections therefore differ by a pure gauge term, $\Gamma^{\prime}-\Gamma=U^{\dagger}{d}U$. In particular, by choosing $U$ to be parameter-independent, this gauge contribution vanishes, ensuring that $\Gamma^{\prime}=\Gamma$., implying that one can set $\Xi^{\lambda}=0$, and thus $\mathbb{A}_{\lambda}$ transforms as an ordinary $U(N)$ gauge potential.
@@ V Example
where $\sigma_{i}$ are Pauli matrices. This two-level Hamiltonian can, for instance, effectively describe topological magnons on a honeycomb lattice [8, 5]. The Hamiltonian in Eq. (17) has also been recently analyzed in Ref. [9] where it was argued that a (2+1)-dimensional electromagnetic field and topological instanton can arise in its biorthonormal eigenspace. Here, by contrast, we show that when the system is treated in the purely quantum regime and analyzed within the present covariant formalism, no synthetic electromagnetic fields emerge in such systems.
The eigenvalues of the Hamiltonian are $E_{\pm}=a_{0}\pm l$, with $l=\sqrt{t^{2}-x^{2}-y^{2}}$, implying that the condition $l\geq 0$ defines an exact pseudo-Hermitian phase where the eigenspectrum is real. Through the hyperbolic parameterization $t=l\cosh\xi$, $x=l\sinh\xi\cos\lambda$, $y=l\sinh\sin\lambda$, the eigenvectors take the form on the reduced 2-dimensional parameter manifold $(\lambda,\xi)$:
@@
For simplicity, we focus on a single isolated energy band with $E_{-}$. By fixing $E_{-}=const$, and exploiting Eqs. (2) and (V), the corresponding conventional left-right Berry connection reads
@@
That is, under the wave-function norm–preserving condition the CBC vanishes identically. As a result, no geometric phase is accumulated when the state undergoes a cyclic evolution, in sharp contrast to the conventional left–right Berry connection in Eq. (19), which in general depends on $\xi$.
@@
(see Appendix C for details). This demonstrates that the CBC is immune to such length-changing transformations of the eigenvectors. Indeed, the affine CBC component in Eq. (12) equals to $\Xi=-\frac{i}{2}\tanh\xi d\xi$ (see Appendix C), which exactly cancels the arising imaginary contribution in the left-right Berry connection in Eq. (20), thus ensuring that the state’s norm is always preserved.
In light of the above, it is also clear that, for this two-level model, no Berry curvature is generated within the CBC formalism, i.e., $F=0$ according to Eq. (15), away from the spectral singularity $l=0$. This is in contrast to the nonzero Berry curvature computed from the left-right Berry connection in Eq. (19)  [^6] Specifically, the Berry curvature computed from the left-right Berry connection in Eq. (19) reads $F=-i(\sinh\xi)\sigma_{y}/2$ [9].. Consequently, in comparison with the results reported in Ref. [9], no $(2+1)$-dimensional electromagnetic fields emerge in the eigenspace of pseudo-Hermitian Hamiltonians in the state-norm-preserving quantum regime  [^7] It should be noted that, while the covariant formalism is essential in the pure quantum regime, the emergent (2+1)D electromagnetic field predicted in Ref. [9] can nevertheless be observed in (semi)classical regime..
These findings can be additionally understood from the perspective of the Hermitizing maps $S$ and $S^{\prime}$ associated with the original and rescaled eigenvectors in Eq. (V). Indeed, as shown in Appendix D, the corresponding eigenvectors of the Hermitian Hamiltonians, $H^{H}=SH^{\rm pH}S^{-1}$ and $H^{\prime H}=S^{\prime}H^{\rm pH}S^{\prime-1}$, namely $|\phi^{H}_{-}\rangle=S|\psi^{R}_{-}\rangle$ and $|\phi^{\prime H}_{-}\rangle=S^{\prime}|\psi^{\prime R}_{-}\rangle$, are equivalent up to an arbitrary phase. Explicitly, one finds
$|\phi^{\prime H}_{-}\rangle=|\phi^{H}_{-}\rangle=\frac{1}{\sqrt{2}}\bigl[1,1\bigr]^{T}.$ (24)
That is, the eigenvector associated with the Hermitian Hamiltonian is fixed across the parameter space, away from the spectral singularity, and the corresponding Berry connection therefore vanishes identically there.

### v2

@@ Abstract
Non-Hermitian systems display spectral and topological phenomena absent in Hermitian physics; yet, their geometric characterization can be hindered by an intrinsic ambiguity rooted in the eigenspace of non-Hermitian Hamiltonians, which becomes especially pronounced in the pure quantum regime. Because left and right eigenvectors are not related by conjugation, their norms are not fixed, giving rise to a biorthogonal ${\rm GL}(N,{\mathbb{C}})$ gauge freedom. Consequently, the standard Berry connection admits four inequivalent definitions depending on how left and right eigenvectors are paired, giving rise to distinct Berry phases and generally complex-valued holonomies. Here we show that these ambiguities and the emergence of complex phases are fully resolved by introducing a covariant-derivative formalism built from the metric tensor of the Hilbert space of the underlying non-Hermitian Hamiltonian. The resulting uniquely defined Berry connection remains real-valued under an arbitrary ${\rm GL}(N,{\mathbb{C}})$ frame change, and transforms as an affine gauge potential, while reducing to the conventional Berry (or Wilczek–Zee) connection in the Hermitian limit. This establishes an unambiguous and gauge-consistent geometric framework for Berry phases, non-Abelian holonomies, and topological invariants in quantum systems described by non-Hermitian Hamiltonians.
@@ I Introduction
Non-Hermitian Hamiltonians naturally emerge in a wide variety of experimental platforms, including photonic systems with engineered gain and loss [32, 14, 31], mechanical and electrical metamaterials with active or nonreciprocal elements [16, 35], acoustic lattices [48, 21], cold-atom systems with controlled dissipation [42, 4], circuit-QED and microwave platforms [33], as well as synthetic quantum matter  [31, 7] to name a few. In these systems, the nontrivial phenomena such as non-Hermitian skin effects [47, 25, 27] and non-Bloch band topology [47, 9, 30] demonstrate that their geometric properties encode unique features not encountered in Hermitian counterparts.
A defining feature of the non-Hermitian Hamiltonians is the biorthogonality of their eigenvectors: right and left eigenstates are independent, and need not be related by Hermitian conjugation [28, 10]. As a consequence, each eigenmode retains an intrinsic ‘normalization’ gauge freedom, namely,
@@
whose transformations under $c_{n}(\boldsymbol{\lambda})$ differ because left and right vectors transform oppositely [15, 37, 41]. As a result, these four connections are generically not gauge equivalent [45]. Moreover, for any parameter-dependent scaling with $|c_{n}|\neq 1$, all ${\cal A}^{\alpha\beta}$ acquire different gauge-dependent imaginary components, rendering the associated Berry phases complex and curvatures non-unique [37]. Specifically, under such rescaling in Eq. (1) with $c_{n}=r_{n}e^{i\phi_{n}}$, the left-right Berry connection for a single-band transforms as
@@
Clearly, for an arbitrary real-valued scaling factor $r_{n}\neq 1$, the Berry connection becomes complex-valued even if the original connection ${\cal A}^{LR}$ is purely real. This implies that one can induce arbitrary geometric exponential damping or amplification of an adiabatically evolving state along open-path trajectories merely by rescaling the right eigenframe, regardless of whether the spectrum of the underlying non-Hermitian Hamiltonian is real or not.
While such complex geometric phases pose no conceptual difficulty in classical and semiclassical dissipative systems, where the evolving state can be regarded as a physical field that may naturally attenuate or amplify along a given trajectory in parameter space [38], and where such geometric effects can even be observed and exploited, as recently demonstrated in Refs. [41, 26], the situation is fundamentally different in the quantum domain. When the wave function encodes probability amplitudes of the system’s possible configurations, rather than a physical field, its norm represents the total probability and must therefore remain fixed under any admissible continuous evolution, whether generated by a Hermitian or a non-Hermitian operator [13, 5]. Although it is common to enforce norm preservation through ad hoc renormalization of the quantum state, this procedure raises concerns about the consistency of such an approach in genuine non-Hermitian quantum settings [39]. These difficulties have motivated recent efforts to formulate the real-valued Berry connection for non-Hermitian systems [45, 46]; however, the applicability of such approaches has so far been restricted to systems with degenerate eigenspaces.
Consequently, this naturally raises the question of whether one can define a unique and real-valued Berry connection for finite-dimensional non-Hermitian Hamiltonians that applies uniformly to both non-degenerate and degenerate eigenspaces and remains real under arbitrary ${\rm GL}(N,\mathbb{C})$ frame transformations. Such a connection, which ensures the geometric norm preservation of the adiabatically evolving states, would also lead to a consistent geometric and topological structure on the eigenvector bundle of the quantum non-Hermitian evolution operators. In this work, we construct such a Berry connection and show that it naturally emerges from the metric associated with the deformed non-Hermitian Hamiltonian Hilbert space.
@@ II Metric tensor and covariant derivative in the Hilbert space of non-Hermitian Hamiltonians
corresponding to, in general, a time-dependent Hermitian Hamiltonian $H^{H}=SHS^{-1}+i(\partial_{t}S)S^{-1}$ [23]. The latter expression coincides with the gauge transformation law for the Hamiltonian viewed as a gauge potential, with the transformation rendering it Hermitian. Indeed, combining Eqs. (5)–(7) one immediately obtains $\langle\phi^{H}|\phi^{H}\rangle=1$, which is a standard definition of the inner product on the Hilbert space of Hermitian Hamiltonians (see also Fig. 1111Note that the Hermitizing map $S$ possesses a unitary gauge freedom and thus is not unique [23]. Indeed, by setting $S\to US$, with some unitary $U$, the metric $\eta$ in Eq. (6) remains invariant. To avoid this ambiguity, one can choose a minimal $S$, whose parameter dependence stems solely from the metric, i.e., $S(\boldsymbol{\lambda})=\sqrt{D(\boldsymbol{\lambda})}W^{\dagger}(\boldsymbol{\lambda})$, such that $\eta(\boldsymbol{\lambda})=WDW^{\dagger}$, with $W\in U(N)$, and $D$ is a positive-definite diagonal matrix. As such, here, we implicitly assume that if there is such a $U$, it is always parameter independent.
@@
where the $S^{-1}\partial_{\lambda}S=\Gamma^{\lambda}$ term serves as the connection component on the $\eta$-deformed Hilbert space, and which obeys the metric-compatible equation for the connection: $\partial_{\lambda}\eta=\eta\Gamma^{\lambda}+\Gamma^{\lambda\dagger}\eta$  [^2] More specifically, $\partial_{\theta}\langle\psi|\psi\rangle_{\eta}=\langle D_{\theta}\psi|\psi\rangle_{\eta}+\langle\psi|D_{\theta}\psi\rangle_{\eta}$=0, implying $\langle D_{\theta}\psi|\psi\rangle_{\eta}=-\langle\psi|D_{\theta}\psi\rangle_{\eta}$. Importantly, this condition does not enforce parallel transport of the state, determined by the expression $D_{\lambda}|\psi\rangle=0$, but merely guarantees norm preservation of a state when it is transported in $\boldsymbol{\lambda}$-space [22]. This situation is fully analogous to the Hermitian case ($\Gamma\to 0$): although the norm of an eigenstate $|\phi^{H}\rangle$ is fixed throughout parameter space $\partial_{\lambda}\langle\phi^{H}|\phi^{H}\rangle=0$, the state itself generally varies with system parameters, i.e., $\partial_{\lambda}|\phi^{H}\rangle\neq 0$. It is precisely this parametric variation that determines the Berry connection..
Furthermore, this connection is curvature-free, as it is generated by a pure-gauge structure (i.e., the Maurer-Cartan form) $S^{-1}dS$, implying $D_{\mu},D_{\nu}]=0 [$29, 22]. Meaning that the Hilbert space, endowed with the metric $\eta$, is intrinsically flat despite the seeming deformation (see Fig. 1).
@@ III Covariant Berry connection
In the Hermitian limit, for a fixed eigenframe, the CBC can be reduced to the standard Berry connection, since $\eta\to I$ and $D_{\lambda}\to\partial_{\lambda}$. Likewise, when the metric constant, that is $\Gamma^{\lambda}=0$, the CBC is simplified to the usual left-right non-Hermitian Berry connection given in Eq. (2). In this sense, the standard non-Hermitian left–right connection ${\cal A}^{LR}$ is appropriate in the genuine quantum regime only when the Hilbert-space metric is fixed (i.e., $\eta=\rm const$). Moreover, via the similarity transformation in Eq. (7), the CBC can also be written as ${\mathbb{A}}^{\lambda}_{mn}=i\langle\phi^{H}_{m}|\partial_{\lambda}|\phi^{H}_{n}\rangle$, which is the conventional Berry connection for a Hermitian system. Thus the CBC in Eq. (9) is manifestly real-valued.
@@
Here, the distortion tensor $\Xi$ is determined by the difference of the metric connections $\Gamma^{\prime\lambda}-\Gamma^{\lambda}$, originating from the non-unitary (positive-definite) part $P$ in the polar decomposition of $T=UP$. In other words, the distortion compensates for the local rescaling of eigenvector norms induced by $P$. In particular, for a purely unitary transformation $T=U$, the metric remains invariant $\eta^{\prime}=\eta$, which makes $(\Gamma^{\prime}-\Gamma)$ a pure gauge  [^3] As noted earlier in the text, for the fixed metric $\eta$, the two connections $\Gamma$ and $\Gamma^{\prime}$ may arise from the two Hermitized maps $S$ and $S^{\prime}=US$, respectively, which differ by a smooth unitary $U$. According to the definition of the connection in Eq. (8), these two connections therefore differ by a pure gauge term, $\Gamma^{\prime}-\Gamma=U^{\dagger}{d}U$. By choosing $U$ to be parameter-independent, this gauge contribution vanishes, ensuring that $\Gamma^{\prime}=\Gamma$., implying that one can always set $\Xi^{\lambda}=0$, and thus $\mathbb{A}_{\lambda}$ transforms as an ordinary $U(N)$ gauge potential.
@@ V Example
where $\sigma_{i}$ are Pauli matrices. This two-level Hamiltonian can, for instance, effectively describe topological magnons on a honeycomb lattice [11, 6]. The same model was recently analyzed in Ref. [12], in the context of emerging electromagnetic fields in the eigenspace of equivalent pseudo-Hermitian Hamiltonians. Here, by contrast, we show that when the system is treated in the purely quantum regime and analyzed within the present covariant formalism, no synthetic electromagnetic fields emerge in such systems.
The eigenvalues of the Hamiltonian are $E_{\pm}=a_{0}\pm l$, with $l=\sqrt{t^{2}-x^{2}-y^{2}}$, implying that the condition $l>0$ defines an exact pseudo-Hermitian phase where the eigenspectrum is real. Through the hyperbolic parameterization $t=l\cosh\xi$, $x=l\sinh\xi\cos\lambda$, $y=l\sinh\xi\sin\lambda$, the eigenvectors take the form on the reduced 2-dimensional parameter manifold $(\lambda,\xi)$:
@@
Following Ref. [12], for simplicity, we focus on a single isolated energy band with $E_{-}$. By fixing $E_{-}=const\neq a_{0}$, and exploiting Eqs. (2) and (V), the corresponding conventional left-right Berry connection reads
@@
That is, under the wave-function norm–preserving condition, the CBC vanishes identically. As a result, no geometric phase is accumulated when the state undergoes a cyclic evolution. This is in contrast to the conventional left–right Berry connection in Eq. (19), which generally depends on $\xi$.
@@
see Appendix C for details. This demonstrates that the CBC is insensitive to such length-changing transformations of the eigenvectors. Indeed, the affine component of the CBC in Eq. (12), $\Xi=-\frac{i}{2}\tanh\xi d\xi$ (see Appendix C), exactly cancels the arising imaginary contribution in the left-right Berry connection in Eq. (20), thus ensuring norm preservation of the quantum state under adiabatic evolution.
In light of the above, it is also clear that, for this two-level model, under the considered constraint $E_{-}=const$, no Berry curvature is generated within the CBC formalism, i.e., $F=0$ according to Eq. (15), in the entire $(\lambda,\xi)$-space. This is in contrast to the nonzero Berry curvature computed from the left-right Berry connection in Eq. (19)  [^6] Specifically, the Berry curvature computed from the conventional left-right Berry connection in Eq. (19) reads $F=-i(\sinh\xi)\sigma_{y}/2$ [12].. Consequently, in comparison with Ref. [12], no ‘fields’ emerge in the eigenspace of such a pseudo-Hermitian system in the quantum regime. We note, however, that the emerging fields, studied in Ref. [12], might still manifest when the Hamiltonian effectively describes (semi)classical systems, where the norm preservation is not essential.
The findings above can be additionally understood from the perspective of the Hermitizing maps $S$ and $S^{\prime}$ associated with the original and rescaled eigenvectors in Eq. (V). Indeed, as shown in Appendix D, the corresponding eigenvectors of the Hermitian Hamiltonians, $H^{H}=SH^{\rm pH}S^{-1}\equiv l\sigma_{x}$, and $H^{\prime H}=S^{\prime}H^{\rm pH}S^{\prime-1}\equiv l\cosh\xi\sigma_{x}$, namely $|\phi^{H}_{-}\rangle=S|\psi^{R}_{-}\rangle$ and $|\phi^{\prime H}_{-}\rangle=S^{\prime}|\psi^{\prime R}_{-}\rangle$, are equivalent up to an arbitrary phase, with the globally smooth $S$ and $S^{\prime}$. Explicitly, one finds
$|\phi^{\prime H}_{-}\rangle\equiv|\phi^{H}_{-}\rangle=\frac{1}{\sqrt{2}}\bigl[1,1\bigr]^{T}.$ (24)
That is, the eigenvectors associated with the Hermitian Hamiltonian are fixed throughout the entire two-dimensional system parameter space $(\xi,\lambda)$, and the corresponding Berry connections, therefore, vanish identically in this region.